OpenStudy (anonymous):

Need help solve with this : (n+1) ^3

OpenStudy (anonymous):

Can you show some steps Please :)

OpenStudy (anonymous):

=(n+1)(n+1)(n+1) you can expand this or use the binomial theorem or the pattern in Pascal's triangle =n^3 + 3n^2 + 3n +1

OpenStudy (anonymous):

is that it :)

OpenStudy (anonymous):

yes - do you know the binomial theorem or are familiar with Pascals triangle?

OpenStudy (anonymous):

no ididnt learn that yet

OpenStudy (anonymous):

(n+1)(n+1)(n+1)= (n^2+2n+1)(n+1) multiply the 1st term of the ( ) into each term of the 2nd ( ) and combine like terms (n+1)(n^2+2n+1) multiply each term in the first ( ) into the 3 terms in the 2nd ( ) and combine like terms. Helpful?

OpenStudy (anonymous):

Thanks